Plot the point given in polar coordinates. (-4, 405°)
Use the given information to find the exact value of the expression. ß lies in quadrant II Find cos (a + cos a = - . a lies i this is a+b
Answer #1

giren that (-4,405°) hell 405° is coterminal to 45° so (-4,45°) = (0,0) * x=rcoso, y=rsino H = -4005(45°) 2 - 2 y = -4sino, =